# Temporal and Semantic Evaluation Metrics for Foundation Models in Post-Hoc Analysis of Robotic Sub-tasks

### v1

@@ Abstract
Recent works in Task and Motion Planning (TAMP) show that training control policies on language-supervised robot trajectories with quality labeled data markedly improves agent task success rates. However, the scarcity of such data presents a significant hurdle to extending these methods to general use cases. To address this concern, we present an automated framework to decompose trajectory data into temporally bounded and natural language-based descriptive sub-tasks by leveraging recent prompting strategies for Foundation Models (FMs) including both Large Language Models (LLMs) and Vision Language Models (VLMs). Our framework provides both time-based and language-based descriptions for lower-level sub-tasks that comprise full trajectories. To rigorously evaluate the quality of our automatic labeling framework, we contribute an algorithm SIMILARITY to produce two novel metrics, temporal similarity and semantic similarity. The metrics measure the temporal alignment and semantic fidelity of language descriptions between two sub-task decompositions, namely an FM sub-task decomposition prediction and a ground-truth sub-task decomposition. We present scores for temporal similarity and semantic similarity above 90%, compared to 30% of a randomized baseline, for multiple robotic environments, demonstrating the effectiveness of our proposed framework. Our results enable building diverse, large-scale, language-supervised datasets for improved robotic TAMP.
@@ I Introduction
Training robotic agents on language-supervised robot trajectory data generally improves their success rates for Task and Motion Planning (TAMP) applications [1, 2]. However, the scarcity of large-scale, language-annotated robot trajectory data, particularly at the granular sub-task level, hinders mass adoption of these methods. The lack of clean, diverse, large-scale and accessible datasets also hampers the creation of Foundation Models (FMs) [3] tailored to robotic use cases [2, 4]. This prevents robotic researchers from reaching an “ImageNet Moment” [5]: a moment which revolutionized the field of computer vision as a result of creating a diverse and vast dataset of labeled images.
Recent works tackle the issue of data scarcity by focusing their efforts on basic hurdles: manually collecting, cleaning, and curating new or existing datasets into massive, open-source repositories [1, 2, 4]. However, efforts to scale-up language annotations, despite their well-known benefits, are often neglected. Prior works that offer automated language annotation methods predominantly generate high-level language instructions for full trajectories [6] or rely on computationally expensive validation methods [7].
@@
Figure 1: Our approach evaluates a Foundation Model’s (FM) ability to temporally and semantically decompose a robot trajectory into a sub-task decomposition. We compare an FM sub-task decomposition, $\hat{\mathcal{S}}$, with a ground-truth, human labeled sub-task decomposition, $\mathcal{S}$, through our core contributions of temporal and semantic alignment metrics.
@@
This paper proposes an alternative framework, shown in Fig. 1, which leverages powerful FMs including Visual Language Models (VLMs) and Large Language Models (LLMs) to automatically generate natural language descriptions and annotations for robot trajectories at sub-task level. In this paper, the VLMs we use include GPT-4V [8], Gemini Pro Vision, [9], and Video-LLaVa [10], and the LLM we use is Gemini Pro. Compared to traditional manual annotation via platforms like MTurk [11], our approach offers significant advantages: scalability and cost-effectiveness. More concretely, a human annotator earning the median wage, \$14.41 per hour [12], would need to annotate full trajectories in less than 30 seconds in order to be as cost effective as our method with GPT-4V. Additionally, our method allows for incorporating rich, nuanced language at a granular sub-task level into robot instructions, improving the capabilities of robotics and Embodied AI agents [13] in understanding a wider range of instructions for Task and Motion Planning. Lastly, our framework is both hardware and simulation environment agnostic.
@@
To ensure the quality of FM-generated annotations, we define two metrics: temporal similarity, $\tau_{k}$, and semantic similarity, $\tau_{\zeta}$. Both metrics compare ground-truth annotations to FM-generated annotations, quantifying their alignment and effectiveness. These metrics allow researchers to assess the quality of the proposed framework when applied to novel tasks. This methodology only requires a test set, i.e, a small sample of labeled data, to evaluate the general performance of the FM-driven framework. Collecting a test set is far less burdensome than collecting a dataset for finetuning. Our experimental results, accumulated over 200 unique robot trajectories with 2400 variations of FM calls, demonstrate the proposed framework produces valid annotations with a $\tau_{k}$ and $\tau_{\zeta}$ both above 90% for contextually informative prompts compared to a randomized baseline near 30%, indicating our method’s effectiveness in adhering to the semantic and temporal realities of the ground-truth.
Our core contributions are:
@@
  • Results of the metrics applied across concrete examples, showcasing the efficacy of FMs in decomposing complex, multi-modal data.
@@ II Related Work
Creating Natural Language Annotations for Robot Trajectories: Prior works generally rely on human annotators to create robot trajectories and their corresponding language descriptions [14, 15, 16, 17, 1, 18, 2]. Our focus is on developing automated methods to generate this data.
There are two works which closely relate to our proposed framework. First, Ha et al. [7] develop a framework that iteratively queries LLMs to break down high-level task instructions into sub-tasks with natural language (NL) descriptions. These subtasks are then grounded in 6-DoF action primitives which create executable trajectories. The trajectory data is vetted via simulation which provides either a success or failure signal. The failure signal initiates a re-querying of the LLM to restart the search for valid trajectory data. We find our method complementary to Ha et al.’s [7] framework, as our method relies on ground-truths generated by a Finite State Machine (FSM) controller, reducing the need for expensive re-queries to the FM. Additionally, our novel similarity metrics can be introduced prior to Ha et al.’s [7] primitive grounding step, further reducing the need for FM re-queries. The scores provided by these metrics can indicate FM response quality without running the full trajectory through a simulation. Moreover, semantic alignment, $\tau_{\zeta}$, provides finer-grain measurement over language quality than task performance metrics, which can miss nuance and noise in the synthetically generated language data. Second, Xiao et al. [6] leverage VLMs to provide automatic NL annotations to robot trajectories; but their methodology can be expensive, requiring roughly 5,600 human-labeled trajectories to fine-tune an off-the-shelf model to produce annotations. By contrast, we use prompting and rely on state-of-the-art FMs’ reasoning and in-context-learning capabilities to produce the annotations. Our approach only leverages human annotations for evaluating the quality of FM-produced annotations via our proposed SIMILARITY algorithm. Lastly, our approach produces annotations at sub-task level whereas Xiao et al. [6] provide high-level descriptions for entire trajectories.
Alternative methods at the intersection of task planning and Explainable AI (XAI) [19] leverage FMs to provide annotations for TAMP. These works are often too fine-grained, relying on the granularity of ROS 2 log files [20]. These works are also often too high-level, providing high-level summaries of trajectories [21] that cannot be used as training data. While explainability is not the primary goal of our work, we recognize this as a future application of our method.
@@
Decomposing Trajectories: Raj et al. [22] treat trajectory decomposition as semantic parsing and are similarly focused on detecting sub-tasks to train a detection algorithm with labeled video clips and language annotation pairs. Prior works in computer vision focus on similar concepts to alleviate manual annotation of video segments, [23], [24], [25]. These methods rely on training detection algorithms rather than prompting multi-modal FMs.
FMs + Task and Motion Planning: Liu et al. [26] combines the strengths of classical planners with modern LLMs to develop long-horizon optimal plans from PDDL queries. Singh et al. [27] prompt LLMs with program-like specifications of the available actions primitives, objects in an environment, and executable, example programs to produce plans. Ahn et al. [28] ground the LLM-generated plans by multiplying each candidate action’s probability under FLAN [29] with the action’s value function, which serves as a proxy for affordance. Huang et al. [30] leverage Chain-of-Thought prompting techniques [31] to produce LLM-based plans. While there is an ever growing literature of FM-based planners [32, 33, 34], our method provides data which supports these applications. Put simply, we perform post-processing of trajectories to develop language-supervised sub-tasks to be used as training data. We do not directly target FM-based planning for TAMP.
@@ III-A Preliminaries and Definitions
A trajectory, $T$, is a complete sequence of robot and non-robot states, $x\in X\subset\mathbb{R}^{p}$, control inputs, $u:X\mapsto U\subset\mathbb{R}^{q}$, and 3 channel images, $\mu\in\mathcal{I}\subset\mathbb{R}^{H\times W\times 3}$, that describes the evolution of robot states and non-robot states over all time steps. Robot states may encompass the x-y-z positions of a robot’s end-effector, whereas non-robot states may include the x-y-z positions of external objects. We assume a trajectory is propagated forward by discrete-time dynamics with step, $k\in[0,K]\subset\mathbb{N}^{+}$, as:
@@
where $f:\mathbb{R}^{p}\times\mathbb{R}^{q}\mapsto\mathbb{R}^{p}$ serves as the dynamics for a robotic agent within a particular environment. The $k$ in $x_{k}$, $u_{k}$, and $\mu_{k}$ denotes the signal at step $k$. We denote the trajectory “length” as the total number of steps, $K$. While $\mu_{k}$ is not directly in Eq. 1, $\mu_{k}$ captures a visual representation of the environment. We consider the control policy to be a high-level TAMP controller, for example a Finite State Machine (FSM) or a Behavior Tree (BT) [35], [36].
@@
$\mathcal{D}=\{(x_{k},u_{k},\mu_{k})\}_{k=0}^{K}.$
We denote $\mathcal{D}^{kxu}\subset\mathcal{D}$ as textual data, which only contains $\{(x_{k},u_{k})\}_{k=0}^{K}$, and denote $\mathcal{D}^{\mu}\subset\mathcal{D}$ as visual data, which only contains $\{\mu_{k}\}_{k=0}^{K}$, so that $\mathcal{D}^{kxu}\cup\mathcal{D}^{\mu}=\mathcal{D}$.
@@ III-B Robotics Manipulator Example
Consider a robotic environment Stack that consists of a manipulator picking up Cube A, stacking Cube A on Cube B, then returning to a home position. Each state at a step, $x_{k}$, consists of the robot’s x-y-z end-effector position and the non-robot states, which here are each cubes’ x-y-z position and the end-effector x-y-z relative position to each cubes’ x-y-z position. The control signal at each step, $u_{k}$, is the desired manipulator joint pose. The image at each step, $\mu_{k}$, is a $256\times 256$ RGB image. The task planner is a Finite State Machine, which carefully sequences sub-tasks together to achieve the goal of stacking Cube A on Cube B.
@@ IV-A FMs for Auto-Decomposition of Robotic Sub-tasks
The core C piece shown in Section V to enable high prediction quality is the “One-Shot” example, $\textbf{C}^{1S}$, shown by IN_CONTEXT_EXAMPLE in Fig. 2. The $\textbf{C}^{1S}$ is actually a snippet of $\mathcal{D}^{kxu}$ (from a hold-out $\mathcal{D}$, $\mathcal{S}$ pair) for each environment to provide only a partial example. When IN_CONTEXT_EXAMPLE is omitted, we consider the C to include $\textbf{C}^{0S}$ rather than $\textbf{C}^{1S}$.
@@ IV-B Similarity Metrics Formulation
Ground-truth $\mathcal{S}$ $\hat{\mathcal{S}}$ using $\mathcal{D}^{kxu}$, $\textbf{C}^{1S}$ $\tau_{k}=0.87$, $\tau_{\zeta}=0.98$ $\hat{\mathcal{S}}$ using $\mathcal{D}^{kxu}$, $\textbf{C}^{0S}$ $\tau_{k}=0.74$, $\tau_{\zeta}=0.51$ Random $\tilde{\mathcal{S}}$ $\tau_{k}=0.31$, $\tau_{\zeta}=0.14$
{(0, 10, “Move to above   xxxxxxx Cube A”), (11, 23, “Move directly down   xxxxxxx to Cube A”), (24, 25, “Grasp Cube A”), (26, 39, “Vertically pick   xxxxxxx up Cube A”), (40, 48, “Align Cube A   xxxxxxx with Cube B”), (49, 54, “Move Cube A   xxxxxxx vertically down   xxxxxxx to Cube B”), (55, 58, “Release Cube A   xxxxxxx onto Cube B”), (59, 62, ‘Return Home”)} {(0, 10, “Move to above   xxxxxxx Cube A”), (11, 23, “Move directly   xxxxxxx down to Cube A”), (24, 25, “Grasp Cube A”), (26, 39, “Vertically pick   xxxxxxx up Cube A”), (40, 54, “Align Cube A   xxxxxxxwith Cube B”), (55, 58, “Release Cube A   xxxxxxx onto Cube B”), (59, 62, “Return Home’)} {(0, 10, “robot arm moving   xxxxxxx towards cube A”), (11, 24, “robot arm adjusting   xxxxxxx position above cube A”), (25, 39, “robot arm descending   xxxxxxx towards cube A”), (40, 54, “robot arm   xxxxxxx lifting cube A”), (55, 62, “robot arm   xxxxxxx moving cube A   xxxxxxx towards cube B”)} {(0, 38, “Sort objects by color”’), (39, 42, “Turn on a light”), (43, 49, “Serve a meal”), (50, 53, “Remove object from wall”), (54, 61, “Place object on a shelf”), (62, 62, “Sort objects by color”)}
TABLE I: Ground-truth $\mathcal{S}$ compared against two $\hat{\mathcal{S}}$ and one $\tilde{\mathcal{S}}$ for a single trajectory in the Stack environment: The second column shows a GPT-4V $\hat{\mathcal{S}}$ using textual data, $\mathcal{D}^{kxu}$, and an in-context example, $\textbf{C}^{1S}$. The third column shows a GPT-4V $\hat{\mathcal{S}}$ using textual data, $\mathcal{D}^{kxu}$, and excludes an in-context example, $\textbf{C}^{0S}$. The last column shows a randomly generated sub-task decomposition, $\tilde{\mathcal{S}}$. The metrics, $\tau_{k}$ and $\tau_{\zeta}$, shown at the top of each column are scores associated with $\texttt{SIMILARITY}(\mathcal{S},\hat{\mathcal{S}})$ or $\texttt{SIMILARITY}(\mathcal{S},\tilde{\mathcal{S}})$. The important note is $\hat{\mathcal{S}}$ using $\textbf{C}^{1S}$ is nearly identical to the ground-truth, while $\hat{\mathcal{S}}$ using $\textbf{C}^{0S}$ has different temporal partitions and overall different vernacular.
@@
After all sub-task intersections are calculated, 10 and 11 calculate $\tau_{k}$ and $\tau_{\zeta}$ to be within their respective ranges.

### v2

@@ Abstract
Recent works in Task and Motion Planning (TAMP) show that training control policies on language-supervised robot trajectories with quality labeled data markedly improves agent task success rates. However, the scarcity of such data presents a significant hurdle to extending these methods to general use cases. To address this concern, we present an automated framework to decompose trajectory data into temporally bounded and natural language-based descriptive sub-tasks by leveraging recent prompting strategies for Foundation Models (FMs) including both Large Language Models (LLMs) and Vision Language Models (VLMs). Our framework provides both time-based and language-based descriptions for lower-level sub-tasks that comprise full trajectories. To rigorously evaluate the quality of our automatic labeling framework, we contribute an algorithm SIMILARITY to produce two novel metrics: temporal similarity and semantic similarity. The metrics measure the temporal alignment and semantic fidelity of language descriptions between two sub-task decompositions, namely an FM sub-task decomposition prediction and a ground-truth sub-task decomposition. We present scores for temporal similarity and semantic similarity above 90%, compared to $\approx$60for human annotations, for multiple, simulated robotic environments, demonstrating the effectiveness of our proposed framework. Our results enable building diverse, large-scale, language-supervised datasets for improved robotic TAMP.
@@ I Introduction
Training robotic agents on language-supervised robot trajectory data generally improves their success rates for Task and Motion Planning (TAMP) applications [1, 2]. However, the scarcity of large-scale, language-annotated robot trajectory data, particularly at the granular sub-task level, hinders mass adoption of these methods. The lack of clean, diverse, large-scale and accessible datasets also hampers the creation of Foundation Models (FMs) [3] tailored to robotic use cases [2, 4]. This prevents robotic researchers from reaching an “ImageNet Moment” [5]: a moment that revolutionized computer vision by creating a vast and diverse labeled dataset.
@@
Figure 1: Our approach evaluates a Foundation Model’s (FM) ability to temporally and semantically decompose a robot trajectory into a sub-task decomposition. We compare an FM sub-task decomposition, $\hat{\mathcal{S}}$, with a ground-truth sub-task decomposition, $\mathcal{S}$, through our core contributions of temporal and semantic alignment metrics. The image above shows how robot trajectory data, $\mathcal{D}$, is post-hoc processed by an FM to compute a predicted sub-task decomposition, $\hat{\mathcal{S}}$, and quantitatively compared to a ground-truth sub-task decomposition, $\mathcal{S}$.
Recent works address data scarcity by manually collecting, cleaning, and curating new or existing datasets into large, open-source repositories [1, 2, 4]. However, efforts to scale language annotations, despite their well-known benefits, are often neglected. Prior works that offer automated language annotation methods predominantly generate high-level language instructions for full trajectories [6] or rely on computationally expensive validation methods [7]. By contrast, we develop rich, nuanced language descriptions at a granular sub-task level towards improving the robotics and Embodied-AI agents [8] capabilities in understanding a wider range of TAMP instructions via refined language supervision.
@@
This paper proposes a sub-task annotation framework, see Fig. 1, which leverages FMs including state-of-the-art Visual Language Models (VLMs) [9, 10, 11] and Large Language Models (LLMs) to automatically generate natural language descriptions and annotations for robot trajectories at sub-task level. Compared to traditional manual annotation via platforms like MTurk [12], our approach offers significant advantages: scalability and cost-effectiveness. For example, a human annotator earning a minimum wage of \$14.41 per hour [13], would need to complete each annotation in 15-30 seconds to match the cost-effectiveness of our GPT-4V method; in our experiments, we found human annotation took between 2-3 minutes.
@@
To evaluate the quality of each annotation produced by our framework, we define two metrics: temporal similarity, $\tau_{k}$, and semantic similarity, $\tau_{\zeta}$. Both metrics are designed to quantify alignment between a ground-truth annotation and FM-generated annotation, allowing researchers to assess the quality of the proposed framework on their own robotics’ trajectories. This methodology only requires a test set (a small sample of labeled data) to evaluate the its performance instead of collecting a dataset for fine-tuning. Our experimental results, accumulated over 200 unique robot trajectories with 2400 variations of FM calls, demonstrate the proposed framework produces valid annotations with a $\tau_{k}$ and $\tau_{\zeta}$ above 90% for contextually informative prompts compared to human-based annotations below 60%, indicating its effectiveness in adhering to the semantic and temporal realities of the ground-truth.
Our core contributions are as follows.
@@
  • Results of the metrics across concrete examples, compared to human annotations, showcasing the efficacy of FMs in decomposing complex, multi-modal data.
@@ II Related Work
Creating Natural Language Annotations for Robot Trajectories: Traditional approaches to generating language descriptions for robot trajectories rely heavily on human-created annotations [14, 15, 16, 17, 1, 18, 2]. In contrast, our work aims to automate this process. To this end, several prior works have developed frameworks that leverage automated techniques for generating task-level descriptions in robotic contexts. Notably, Ha et al. [7] propose a framework that iteratively queries LLMs to break down high-level task instructions into sub-tasks with natural language (NL) descriptions. Our method could extend this approach by using a Finite-State Machine (FSM) controller to validate annotations, reducing the need for costly FM re-queries. Additionally, our semantic alignment metric, $\tau_{\zeta}$, enables a more detailed assessment of language quality compared to task performance metrics, which can overlook subtle nuances and noise in synthetic language data. Other FM-based methods automate trajectory labeling through object detection and heuristics [19] or require extensive fine-tuning on human-labeled data [6]. By contrast, our approach relies on prompt-based in-context learning for FM reasoning, using human annotations solely to evaluate the SIMILARITY algorithm. Our method uniquely generates sub-task-level, temporally grounded annotations rather than descriptions for entire trajectories, enabling more precise alignment across trajectory steps.
Semantic Preservation in Robotic Task Embeddings: Other work, such as Cao et al. [20], preserves the semantic structure of behavior trees (BTs) in robotic tasks by embedding the BT hierarchy and sequence into representations using language embedding models like USE [21] and GloVe [22]. Unlike our approach, which focuses on assessing annotation quality rather than behavioral coordination, Cao’s method organizes descriptors within the BT structure. Their approach encodes sequential information through unitless exponential decay, lacking explicit temporal grounding unlike our method. However, we both employ cosine-based similarity on USE embeddings for validating semantics in robotic contexts.
Alternative approaches at the intersection of task planning and Explainable AI (XAI) [23] use FMs to provide annotations for TAMP or generate summaries describing robot behavior [24]. Some of these methods require highly detailed data, such as ROS2 log files [25], which limit their general applicability. Explainability is not the primary focus of our work, but is a potential future application of our method.
@@
Decomposing Trajectories: [26] treat trajectory decomposition as semantic parsing and are similarly focused on detecting sub-tasks to train a detection algorithm with labeled video clips and language annotation pairs. Prior works in computer vision focus on similar concepts to alleviate manual annotation of video segments, [27], [28], [29] which rely on training detection algorithms rather than prompting multi-modal FMs.
Annotation Quality Assurance Metrics: Previous works have primarily evaluated the quality of labeled data using binary success rates on downstream tasks. However, this binary approach overlooks important factors in data generation, such as coverage and diversity, which are critical when developing synthetic datasets. Assessing these factors can also highlight potential shortcomings, such as class imbalance and noise, e.g., temporally unaligned language instructions. Therefore, instead of relying on these binary metrics, we compare the task decomposition performance of our VLM and LLM models with that of human labelers.
@@ III-A Preliminaries and Definitions
A trajectory, $T$, is a complete sequence of robot and non-robot states, $x\in X\subset\mathbb{R}^{p}$, control inputs, $u:X\mapsto U\subset\mathbb{R}^{q}$, and 3 channel images, $\mu\in\mathcal{I}\subset\mathbb{R}^{H\times W\times 3}$, that describes the evolution of robot states and non-robot states over all time steps. Robot states may encompass the robot’s end-effector pose, whereas non-robot states may include the external objects pose. We assume a trajectory is propagated forward by discrete-time dynamics with step, $k\in[0,K]\subset\mathbb{N}^{+}$, as:
@@
where $f:\mathbb{R}^{p}\times\mathbb{R}^{q}\mapsto\mathbb{R}^{p}$ serves as the dynamics for a robotic agent within a particular environment. The $k$ in $x_{k}$, $u_{k}$, and $\mu_{k}$ denotes the signal at step $k$. We denote the trajectory “length” as the total number of steps, $K$. While $\mu_{k}$ is not directly in Eq. 1, $\mu_{k}$ captures a visual representation of the environment. We consider the control policy to be a high-level TAMP controller, for example a FSM or a Behavior Tree (BT) [30], [31].
@@
$\mathcal{D}=\{(k,x_{k},u_{k},\mu_{k})\}_{k=0}^{K}.$
We denote $\mathcal{D}^{kxu}\subset\mathcal{D}$ as textual data, which only contains $\{(k,x_{k},u_{k})\}_{k=0}^{K}$, and denote $\mathcal{D}^{\mu}\subset\mathcal{D}$ as visual data, which only contains $\{\mu_{k}\}_{k=0}^{K}$, so that $\mathcal{D}^{kxu}\cup\mathcal{D}^{\mu}=\mathcal{D}$.
@@ III-B Robotics Manipulator Example
Consider a robotic environment Stack that consists of a manipulator picking up Cube A, stacking Cube A on Cube B, then returning to a home position. Each state at a step, $x_{k}$, consists of the robot’s end-effector pose and the non-robot poses, which are each cubes’ pose and the end-effector pose relative position to each cubes’ pose. The control signal at each step, $u_{k}$, is the desired manipulator joint pose. The image at each step, $\mu_{k}$, is a $256\times 256$ RGB image. The task planner is an FSM, which carefully sequences sub-tasks together to achieve the goal of stacking Cube A on Cube B.
@@ IV-A FMs for Auto-Decomposition of Robotic Sub-tasks
The core C piece shown in Section V to enable high prediction quality is the “One-Shot” example, $\textbf{C}^{1S}$, shown by IN_CONTEXT_EXAMPLE in Fig. 2. The $\textbf{C}^{1S}$ is actually a snippet of $\mathcal{D}^{kxu}$ (from a hold-out $\mathcal{D}$, $\mathcal{S}$ pair) for each environment to provide only a partial example. When IN_CONTEXT_EXAMPLE is omitted, we consider the C to include $\textbf{C}^{0S}$ rather than $\textbf{C}^{1S}$. That is, $\textbf{C}^{0S}$ does not provide any in-context examples in the P.
@@ IV-B Similarity Metrics Formulation
Ground-truth $\mathcal{S}$ $\hat{\mathcal{S}}$ using $\mathcal{D}^{kxu}$, $\textbf{C}^{1S}$ $\tau_{k}=0.87$, $\tau_{\zeta}=0.98$ cost = \$0.12 $\hat{\mathcal{S}}$ using $\mathcal{D}^{kxu}$, $\textbf{C}^{0S}$ $\tau_{k}=0.74$, $\tau_{\zeta}=0.51$ cost = \$0.10 $\tilde{\mathcal{S}}$ using $\mathcal{D}^{\mu}$ $\tau_{k}=0.45$, $\tau_{\zeta}=0.22$ time spent = 133 seconds cost $\approx$ \$0.53
{(0, 10, “Move to above   xxxxxxx Cube A”), (11, 23, “Move directly down   xxxxxxx to Cube A”), (24, 25, “Grasp Cube A”), (26, 39, “Vertically pick   xxxxxxx up Cube A”), (40, 48, “Align Cube A   xxxxxxx with Cube B”), (49, 54, “Move Cube A   xxxxxxx vertically down   xxxxxxx to Cube B”), (55, 58, “Release Cube A   xxxxxxx onto Cube B”), (59, 62, ‘Return Home”)} {(0, 10, “Move to above   xxxxxxx Cube A”), (11, 23, “Move directly   xxxxxxx down to Cube A”), (24, 25, “Grasp Cube A”), (26, 39, “Vertically pick   xxxxxxx up Cube A”), (40, 54, “Align Cube A   xxxxxxxwith Cube B”), (55, 58, “Release Cube A   xxxxxxx onto Cube B”), (59, 62, “Return Home’)} {(0, 10, “robot arm moving   xxxxxxx towards cube A”), (11, 24, “robot arm adjusting   xxxxxxx position above cube A”), (25, 39, “robot arm descending   xxxxxxx towards cube A”), (40, 54, “robot arm   xxxxxxx lifting cube A”), (55, 62, “robot arm   xxxxxxx moving cube A   xxxxxxx towards cube B”)} {(0, 24, “robot arm moves   xxxxxxx over red block”’), (25, 27, “robot arm grips   xxxxxxx red block”), (28, 40 “robot arm lifts   xxxxxxx red block”), (41, 56, “robot arm moves   xxxxxxx red block over   xxxxxxx to the top of   xxxxxxx the green block”), (57, 59, “robot arm releases   xxxxxxx grip of red block’), (60, 62, “robot arm stows away”)}
TABLE I: Ground-truth $\mathcal{S}$ compared against two $\hat{\mathcal{S}}$ and one $\tilde{\mathcal{S}}$ for a single trajectory in the Stack environment: The second column shows a GPT-4V $\hat{\mathcal{S}}$ using textual data, $\mathcal{D}^{kxu}$, and an in-context example, $\textbf{C}^{1S}$. The third column shows a GPT-4V $\hat{\mathcal{S}}$ using textual data, $\mathcal{D}^{kxu}$, and excludes an in-context example, $\textbf{C}^{0S}$. The last column shows a human annotated sub-task decomposition, $\tilde{\mathcal{S}}$. The metrics, $\tau_{k}$ and $\tau_{\zeta}$, shown at the top of each column are scores associated with $\texttt{SIMILARITY}(\mathcal{S},\hat{\mathcal{S}})$ or $\texttt{SIMILARITY}(\mathcal{S},\tilde{\mathcal{S}})$. The key insight is $\hat{\mathcal{S}}$ using $\textbf{C}^{1S}$ is nearly identical to the ground-truth, while $\hat{\mathcal{S}}$ using $\textbf{C}^{0S}$ has different temporal partitions and overall different vernacular. Additionally, the two $\hat{\mathcal{S}}$ are nearly 1/5 of the cost compared to $\tilde{\mathcal{S}}$, and have higher $\tau_{k}$ and $\tau_{\zeta}$ metrics.
@@
The interval weight scales sub-task significance within the trajectory, preventing short overlaps from skewing similarity scores. Unlike IoU, it normalizes across trajectory lengths for consistency.
After all sub-task intersections are calculated, 10 and 11 calculate $\tau_{k}$ and $\tau_{\zeta}$ to be within their respective ranges, i.e. $\tau_{k}\in[0,1]$ and $\tau_{\zeta}\in[-1,1]$.